OpenStudy (anonymous):

help me please??:/ What is the vertex form of the equation? y = x^2 – 6x + 8

OpenStudy (campbell_st):

well you need to complete the square in x \[y = (x^2 - 6x + [\frac{6}{2}]^2) + 8 - (\frac{6}{2})^2\] factor the perfect square and you'll have the vertex form

OpenStudy (anonymous):

can u help me do it?

OpenStudy (campbell_st):

well what is (6/2)^2 = ..?

OpenStudy (anonymous):

9?

OpenStudy (campbell_st):

yep so you have added 9 to complete the square and then subtracted 9 for 8 so you have \[y=(x^2 -6x +9) + 8 - 9\] factor the brackets.... its a perfect square... then 8 - 9 = -1 so you will have \[y = (x - ?)^2 - 1\] whats the value of ?

OpenStudy (campbell_st):

the vertex form is \[y = a(x -h)^2 + k\] (h, k) is the vertex of the parabola. your final factored equation should look like it hope this helps

OpenStudy (anonymous):

how do i find the one thats missing in y = x - ?)^2 -1?

OpenStudy (campbell_st):

well you need to factor the quadratic \[x^2 -6x + 9\] what factors of 9 add to 6.. that will be ?

OpenStudy (anonymous):

a factor of 9 is 3, and 3+3 = 6..?

OpenStudy (campbell_st):

yep... so thats the ? its \[y = (x -3)^2 -1 \] thats the vertex form

OpenStudy (anonymous):

and thats the answer?

OpenStudy (campbell_st):

thats it

OpenStudy (anonymous):

thank you soo much!!:D